# Quasi-Solid and Supersolid from Quasiperiodic Long-Range Interactions

### v1

@@ Abstract
Both fits reproduce the data well within Monte Carlo uncertainties. However, the near-flatness of the power-law decay and the quantitative agreement with the constant-like form strongly suggest that $S(q^{*})$ saturates to a nonzero value as $L\to\infty$. This is consistent with theoretical expectations of finite Bragg-like peaks in Fibonacci-modulated systems, and supports the existence of robust quasiperiodic density modulations. Taken together, the results indicate that the quasi-supersolid phase remains stable in the thermodynamic limit.
@@
Specifically, the dominant peak at $q^{*}=2.398$ corresponds to the primary Fourier component of the Fibonacci modulation at $q^{*}=2\pi F_{k-2}/F_{k}=2\pi\cdot 89/233$, where $F_{k}$ denotes the $k$-th Fibonacci number and $L=233=F_{13}$ in our case. Its mirror-symmetric counterpart appears at $q=3.882\approx 2\pi-q^{*}$, consistent with the intrinsic inversion symmetry of the modulation. In addition, two weaker peaks emerge at $q^{*}=1.483$ and $4.800$, which correspond to higher-order Fourier components. The former arises from the ratio $2\pi\cdot F_{k-3}/F_{k}=2\pi\cdot 55/233$, and the latter is its mirror image $2\pi-q^{*}$. These subleading peaks reflect the self-similar, hierarchical structure of the Fibonacci modulation and indicate the presence of multiple rational approximants of the golden ratio $\alpha=(1+\sqrt{5})/2$ encoded in the density response. Together, the appearance of these four peaks confirms the quasiperiodic nature of the emergent order and matches expectations from the mathematical theory of aperiodic diffraction.
@@
Together with the finite-size scaling results discussed before, these large-system data confirm that the quasi-supersolid phase is not a finite-size artifact, but a genuine many-body state stabilized by the interplay of long-range interactions and quantum fluctuation. While the filling $\langle n\rangle=0.694$ is slightly off the Fibonacci ratio $\langle n\rangle=0.618\approx 1/\alpha$, the persistence of both superfluidity and long-range order underscores the robustness of quasi-supersolid phase across a broad parameter regime.

### v2

@@ Abstract
Both fits reproduce the data well within Monte Carlo uncertainties. The near-flatness of the power-law decay and the quantitative agreement with the constant-like form strongly suggest that $S(q^{*})$ saturates to a nonzero value as $L\to\infty$. This is consistent with theoretical expectations of finite Bragg-like peaks in Fibonacci-modulated systems, and supports the existence of robust quasiperiodic density modulations. Taken together, the results indicate that the quasi-supersolid phase remains stable in the thermodynamic limit.
@@
Specifically, the dominant peak at $q^{*}=2.398$ corresponds to the primary Fourier component of the Fibonacci modulation at $q^{*}=2\pi F_{k-2}/F_{k}=2\pi\cdot 89/233$, where $F_{k}$ denotes the $k$-th Fibonacci number and $L=233=F_{13}$ in our case. Its mirror-symmetric counterpart appears at $q=3.882\approx 2\pi-q^{*}$, consistent with the intrinsic inversion symmetry of the modulation. In addition, two weaker peaks emerge at $q^{*}=1.483$ and $4.800$, which correspond to higher-order Fourier components. The former arises from the ratio $2\pi\cdot F_{k-3}/F_{k}=2\pi\cdot 55/233$, and the latter is its mirror image $2\pi-q^{*}$. These subleading peaks reflect the self-similar, hierarchical structure of the Fibonacci modulation and indicate the presence of multiple rational approximants of the inverse golden ratio $\alpha=(1-\sqrt{5})/2$ encoded in the density response. Together, the appearance of these four peaks confirms the quasiperiodic nature of the emergent order and matches expectations from the mathematical theory of aperiodic diffraction.
@@
Together with the finite-size scaling results discussed before, these large-system data confirm that the quasi-supersolid phase is not a finite-size artifact, but a genuine many-body state stabilized by the interplay of long-range interactions and quantum fluctuation. While the filling $\langle n\rangle=0.694$ is slightly off the Fibonacci ratio $\langle n\rangle=0.618\approx\alpha$, the persistence of both superfluidity and long-range order underscores the robustness of quasi-supersolid phase across a broad parameter regime.